OpenStudy (kichi):

Calculate the magnitude and direction of the force of the moon pulling on you. Could you also measure this force?

OpenStudy (kichi):

will be fan

OpenStudy (hoot_hoot):

Universal Gravitation Equation: F = GMm/R2 where F is the force of attraction between two objects in newtons (N) G is the Universal Gravitational Constant = 6.674*10−11 N-m2/kg2 M and m are the masses of the two objects in kilograms (kg) R is the separation in meters (m) between the objects, as measured from their centers of mass Found this at: http://www.school-for-champions.com/science/gravitation_force_objects.htm Hope it helps.

OpenStudy (irishboy123):

clarifying hoot's eqn: \( F = \dfrac{GMm}{R^2}\) And here's a little discussion about using a pendulum as a really simple way to measure the effect on earth http://astronomy.stackexchange.com/questions/11616/can-i-measure-the-moons-gravity you could try it a bit differently - using the idea that the period of a pendulum is gravity dependent \(T = 2 \pi \sqrt {\dfrac{l}{g}}\) and so when the moon is over head net gravity is reduced.... and increased when it's on the other wide of the world pulling down on you. dunno whether the numbers would make any appreciable difference in T but you could runs some simple calculations to test that..... as we already have all the data we need to cheat :-)

OpenStudy (osprey):

Had I a space rocket in my back pocket, then I MIGHT take it out to the point between the earth and the moon where, it seems, that the two grav pulls "cancel". I think that it's much liked by NASA, which may be where I first heard about it, as in someone SAYING it. Luckily for all concerned, I don't have a space rocket ... nor the money to .... So, it's over to you messrs Gagarin et al. It's also just occurred to me that the force posted about is strong enough, i think, to affect tidal waters on earth/sea, and MAY be able to influence bits of the human body (lunatics = lunar tics), but I wouldn't spend much more than typing time on the last one. bon chance et bon voyage http://perendis.webs.com